For what values of 𝑐 does the curve 𝑓(π‘₯) = 2π‘₯3 + 𝑐π‘₯2 + 2π‘₯ have maximum and minimum






points?

1
Expert's answer
2022-04-27T14:46:01-0400
"f'(x)=6x^2+2cx+2"

Find the critical number(s)


"f'(x)=0=>6x^2+2cx+2=0"

"3x^2+cx+1=0""x=\dfrac{-c\pm\sqrt{c^2-12}}{6}"

We consider "x\in \R"

"c^2-12\ge0=>c\le-\sqrt{12}\ or\ c\ge\sqrt{12}"

The curve "\ud835\udc53(\ud835\udc65) = 2\ud835\udc65^3 + \ud835\udc50\ud835\udc65^2 + 2\ud835\udc65" have maximum and minimum for "c\in(-\infin, -\sqrt{12})\cup (\sqrt{12}, \infin)."